MENU C W - WT% Cg - 89.0 Then - 4.0 | Creater 7 g -97.0 0.614 1.96 1000 than - Tor 100
Im trying to learn how to do equivalent ratios but i dont understand how do it. if i have these three chemical components in a solution how do I calculate the equivalent ratio
Answer #1

Equivalent ratios are calculated by multiplying or dividing the numbers of the ratio, by any same number. For eg. For equivalent ratio of 5:20→

(5*2)/(20*2)=10/40 and

(5/5)/(20/5)=1/4

Therefore ,10/40 and 1/4 are equivalent ratios of 5/20. In your problem, if you have to calculate the equivalent ratio of 4/7

(4*2)/(7*2)= 8/14

Similarly, for 0.61/1.96

(0.6*2)/(1.96*2)= 1.2/3.92

Are the corresponding equivalent ratios